find the cost of painting the all round surface of a box with dimensions 60 cm long,40 cm wide,30 cm high at 50 paise per 20 cm square
Answer by mananth(16946)

60 cm long,40 cm wide,30 cm high at 50 paise per 20 cm squareSurfaxe area of box = 2lb+2wb+2wh
=2*60*40+2*40*30+1*30*60=9000 sq. cm
20 sq.cm Re 0.5
9000 sq.m ------x
x=9000*0.5/20
= Rs. 225
m.ananth@hotmail.ca